# Audit Cards: Contextualizing AI Evaluations

### v1

@@ 4 What do existing evaluation reports include?
Methodology: To study the thoroughness of existing evaluation reports, we examined 20 reports produced between 2023 and 2025. We selected these 20 due to their focus on frontier systems, ensuring to include reports from both developers and third party organizations. We manually annotated each report using the 3 principles and 6 features from our audit card template (Section 3). For each audit card component, we scored reports on a 0-2 scale for comprehensively reporting that component (2), minimally reporting that component (1), not reporting on this component at all (0). We assessed reporting thoroughness only, not execution quality, making no judgments about how well evaluations were executed. This means, for example, transparently reporting that there was a lack of any review/feedback process and justifying this would receive a full score of 2. We summarize the findings in Table 2.
[…]
@@
While all evaluation reports provide at least a general overview of procedures, the level of detail differs. We find that audit reports consistently offer information on evaluation scope (average score 1.25) and procedures (1.15) used. However, contextual details are much less consistently reported (average 0.86, $\sigma=0.65$) both across organizations and sometimes within the same organization. For example, Google DeepMind released four documents with significantly varying levels of contextual detail (average scores of 1.00, 0.67, 0.56, and 1.44 in chronological order). Only some reports (e.g., Meinke et al. ; Gemma Team et al. (2024); Üstün et al. (2024)) offered detailed information on the scaffolding used to evaluate the system.
Reports inconsistently address assumptions, limitations, and auditor information. While justifications and limitations are more commonly acknowledged (18 of 20) than assumptions (12 of 20), they are typically discussed at a high level rather than relating to specific methodological choices. Ten reports discuss limitations minimally, while eight do so comprehensively. However, there are notable qualitative differences even among more comprehensive discussions. For example, Üstün et al. (2024) discuss limitations in prompting techniques, language transferability, evaluation tool chaining, and reproducibility in significant detail, whereas the OpenAI system card (OpenAI, 2024) only consider limitations of what was being evaluated (e.g. classified information or restricted data) rather than substantial limitations of the procedures. Regarding auditor information, 12 of 20 reports include auditor expertise details, but rarely their background (3 of 20), making an assessment of potential biases in the audit harder.
Reports rarely disclose information about evaluation integrity, resources, review processes, and obsolescence criteria. Only 3 of 20 reports detail organization and auditor selection, conflicts of interest, or contractual arrangements governing the evaluation process. Even in these cases, the report tended to be about the contractual details of human baselines that are part of the evaluation rather than the core auditing team itself. Only 7 of 20 reports specify which resources auditors had access to, including model access level, computational resources, and time constraints. Similarly, only six reported on some form of review and 9 on feedback mechanisms (Meinke et al., ; METR, 2024a; OpenAI, 2024; UK AISI & US AISI, 2024a). Four report whether they underwent any form of peer review or quality assurance process before publication (e.g. Gemini Team et al., 2024; UK AISI & US AISI, 2024b). Five specified an option for public feedback and corrections. Zero reports explicitly present obsolescence criteria.
@@ 5 What guidance do current governance frameworks offer?
Methodology: We analyzed key governance frameworks issued by institutions that are considered influential regarding technical evaluations for risks from state-of-the-art models. We selected these frameworks based on jurisdiction, institutional authority, and direct influence on model evaluations (see Appendix D for details). This included official regulations (EU AI Act, US Executive Order), voluntary industry standards (NIST AI Risk Management Framework), and policies from major AI developers. We scored each framework on whether it either explicitly required or recommended reporting of each feature in our audit card template, as shown in Table 3.
[…]
Table 3: What guidance do existing governance frameworks provide for reporting? We give binary scores of 1 (requires or recommends explicit reporting of this feature) or 0 (may mention issue’s importance but not specifically its disclosure or reporting). For company policies, it is scored by whether it is disclosed or not. Citations and details for each framework are in Appendix D.
Existing frameworks offer limited guidance on audit reporting. We find existing frameworks consistently require that models are assessed but do not consistently specify when, by whom, for what, and how. This is somewhat expected, as governance frameworks for emerging technologies tend to crystallize gradually (Linkov et al., 2018). High-level requirements in governance documents are often supplemented with more specific forms of soft law such as codes of practice, guidance, and industry norms. Notably, these frameworks rarely require evaluation reports to disclose matters key to the integrity of the process (only 8 of 21) and the resources and access given to auditors (only 2 of 21).
@@ 6 Challenges with auditing according to expert stakeholders
Methodology: To more thoroughly understand audit reporting and perceived gaps between current and best practices, we interviewed 10 experts. We reached out to these experts based on their expertise and familiarity across the evaluation ecosystem. The 10 experts come from a diverse range of private companies, non-profit organizations, and government bodies. Interviewees include: evaluation designers (P1), evaluation developers (P3, P4, P5), evaluation report writers (P1, P3, P5), evaluation-related policy researchers (P6, P7), and those supporting evaluation delivery and standards development (P8, P9). To allow for candid discussions, we committed to working with the level of anonymity that was comfortable for experts. See Section F.1 and Section F.2 for interviewees’ backgrounds and interview questions.
@@ 7 Discussion and recommendations
There are no established standards around how to report the context of audits. The present lack of standards leads to incompleteness and inconsistency in audit reporting practices. Certain contextual aspects, such as the integrity of auditing processes, are particularly neglected. Lacking transparency around audit methodology and context creates a significant barrier to accountability in the AI ecosystem, where trust in evaluations is essential for informed decision-making by regulators, users, and the public.
Audit cards are a practical solution to reporting that accounts for an audit’s real-world context and constraints. The limitations and inconsistencies in current reporting can limit the ability of evaluations to serve their intended role. As the field grows, having standards for reporting is needed to ensure trust between developers, auditors, regulators, and the public. By offering a checklist for reporting, audit cards are designed to capture essential contextual information and allow for a meaningful interpretation of results.
Audit cards remain one approach among several to increase transparency between developers, users, and other stakeholders. Audit cards are designed to offer a structured checklist applicable to all types of assessment (see Section C.2), but should be viewed as an instrument in a larger AI accountability toolkit. The key context they offer can be helpful to other existing mechanisms, including model and system cards, datasheets, and benchmark details (Mitchell et al., 2019; Gebru et al., 2021; Reuel et al., 2024b). Further research may investigate how these cards can enhance and supplement other transparency methods.
@@
  • Limited sample size and reporting details. To enable consistent high-level analysis of the current ecosystem, we used simple 2 or 3 point scoring methods. However, this comes at the expense of granular details and nuance, differentiating between reports only to a limited degree. This is particularly pronounced for our analysis of governance frameworks, where we use binary scores. Furthermore, our analysis, while extensive (25 academic papers, 20 evaluation reports, and 21 governance frameworks), is not exhaustive.
@@ Acknowledgments
This research was supported by the ML Alignment & Theory Scholars (MATS) Program, which provided funding for Leon Staufer and Mick Yang through research stipends. We also thank MATS and our research manager Juan Gil for his organizational assistance and research support.
@@ Appendix A Disambiguating auditing and the evaluations process
AI audits, for the purpose of our analysis, refer to formalized evaluation processes of AI models or systems. By formalized, we mean that it is performed systematically and by an organization, i.e., not just being performed ad hoc by individuals. This kind of evaluation is a model audit, as opposed to a governance audit or an application audit (Mökander et al., 2024).
The evaluation is the underlying assessment run on the model during the audit. We account for the fact that evaluations can occur internally or externally, and pre- or post-deployment. Evaluations can assess capabilities, propensities, or risks. Evaluations can be in many forms, such as benchmarks, red-teaming, human uplift studies, and user studies. Our approach considers all these variants, as the contextual features we have identified are relevant to all. It is possible, that in some of these contexts certain features will be more important and should therefore be reported in more detail. We explore the importance of certain features further in Section 7 and Section C.1.
@@
When speaking of evaluation reports, we refer to internal or external documents including evaluation results (e.g., a model’s results on a benchmark). Examples of evaluation reports include: (OpenAI, 2024; Meinke et al., ) as well as private evaluation reports produced by third-parties for AI companies and internal reports created by AI companies themselves.
@@ Appendix B Literature review
We reviewed past work on “AI auditing”, “AI evaluations”, “best practices”, “science of evaluations”, “evaluation context” and “reporting”. We then assessed each paper for relevance based off the titles. This only yielded eight papers. We thus expanded to adjacent fields such as transparency and AI auditing, technical evaluation design, and sociotechnical approaches. These fields integrate technical performance metrics with societal assessments and considerations of how technical performance interacts with real-world social, cultural, and institutional factors. We scored recommendations within each paper as central arguments (2), minor arguments (1), and no mention (0). Through four iterations, we refined which aspects were most salient, focusing particularly on aspects that were actionable for auditors. We then categorized and prioritized these aspects based on input from experts. The mapping of aspects to features is shown in Section B.2. The final audit card is depicted in Figure 1 and Appendix C.
B.1 List of papers
@@
    Towards Publicly Accountable Frontier LLMs (Anderljung et al., 2023)
@@
    Science of Evals (Apollo Research, 2024)
@@
    Declare and Justify: Explicit assumptions in AI evaluations are necessary for effective regulation (Barnett & Thiergart, 2024a)
@@
    What AI evaluations for preventing catastrophic risks can and cannot do (Barnett & Thiergart, 2024b)
@@
    AI auditing: The Broken Bus on the Road to AI Accountability (Birhane et al., 2024)
@@
    Structured Access for Third-Party Research (Bucknall & Trager, 2023)
@@
    Rethink reporting of evaluation results in AI (Burnell et al., 2023)
@@
    Black-Box Access is Insufficient for Rigorous AI Audits (Casper et al., 2024)
@@
    A Survey on Evaluation of Large Language Models (Chang et al., 2023)
@@
    Who Audits the Auditors? Recommendations from a field scan of the algorithmic auditing ecosystem (Costanza-Chock et al., 2022)
@@
    Hard choices in artificial intelligence (Dobbe et al., 2021)
@@
    Dimensions of Generative AI Evaluation Design (Dow et al., 2024)
@@
    Can We Trust AI Benchmarks? (Eriksson et al., 2025)
@@
    Datasheets for Datasets (Gebru et al., 2021)
@@
    Devising ML Metrics (Hendrycks & Woodside, 2024)
@@
    Holistic Evaluation of Language Models (Liang et al., 2023)
@@
    Task Development Guide (METR, b)
@@
    Auditing Large Language Models: A Three-Layered Approach (Mökander et al., 2024)
@@
    Reasons to Doubt the Impact of AI Risk Evaluations (Mukobi, 2024)
@@
    Towards AI Accountability Infrastructure (Ojewale et al., 2024)
@@
    Closing the AI Accountability Gap: Defining an End-to-End Framework for Internal Algorithmic Auditing (Raji et al., 2020)
@@
    BetterBench (Reuel et al., 2024b)
@@
    Fairness and Abstraction in Sociotechnical Systems (Selbst et al., 2019)
@@
    Model evaluation for extreme risks (Shevlane et al., 2023)
@@
    Toward an evaluation science for generative AI systems (Weidinger et al., 2025)
[…]
Table 4: Annotations for all 25 papers (see Section B.1) across the 24 initial contextual aspects (see Section B.2). We score each paper’s discussion of these as 2 (central argument), 1 (minor argument), and 0 (no mention).
B.2 Mapping of initial aspects
[…]

### v2

@@ 4 What do existing evaluation reports include?
Methodology: To study the thoroughness of existing evaluation reports, we examined 24 reports produced between 2023 and 2025. We selected these 24 due to their focus on frontier systems, ensuring to include reports from both developers and third party organizations. We manually annotated each report using the 3 principles and 6 features from our audit card template (Section 3). For each audit card component, we scored reports on a 0-2 scale for comprehensively reporting that component (2), minimally reporting that component (1), not reporting on this component at all (0). We assessed reporting thoroughness only, not execution quality, making no judgments about how well evaluations were executed. This means, for example, transparently reporting that there was a lack of any review/feedback process and justifying this would receive a full score of 2. We summarize the findings in Table 2.
Report    Release date Justif. Assum. Limit. Who What How Access Integrity Review Average
Average 1.17 0.67 1.42 0.75 1.38 1.29 0.50 0.17 1.13 0.94
Mistral 7B 10/23 0 0 0 0 1 1 0 0 0 0.22
Apollo 01/24 2 1 2 0 2 2 0 0 2 1.22
Cohere Aya 02/24 2 1 2 2 2 1 0 1 1 1.33
UK AISI evals 05/24 1 0 1 0 1 1 0 0 2 0.67
Gemini paper 06/24 1 1 1 1 1 1 1 1 1 1.00
Llama 3 paper 07/24 1 1 2 2 1 2 0 0 1 1.11
Llama 3 card 07/24 1 1 1 1 1 1 0 0 1 0.78
METR o1 09/24 1 0 1 1 1 1 1 0 1 0.78
Claude 3 card 10/24 1 1 1 1 1 1 0 0 1 0.78
METR Claude 3.5 10/24 1 0 2 1 1 1 0 0 1 0.78
Gemma 2 paper 10/24 1 0 1 0 2 1 0 0 1 0.67
AISI Sonnet 3.5 11/24 2 1 2 1 2 1 1 0 1 1.22
OpenAI o1 card 12/24 1 1 1 1 1 1 1 0 1 0.89
Gemma 2 card 12/24 1 0 1 0 1 1 0 0 1 0.56
Gemini 1.5 paper 12/24 1 1 2 2 1 2 1 1 2 1.44
AISI o1 12/24 2 1 2 1 2 1 1 0 1 1.22
METR update 01/25 1 1 2 0 1 1 0 0 2 0.89
DeepSeek R1 01/25 0 0 0 0 1 1 0 0 0 0.22
o3 mini card 01/25 1 1 1 1 1 1 1 0 1 0.89
DeepSeek V3 02/25 1 0 1 0 1 1 0 0 1 0.56
METR o3 and o4-mini 04/25 2 1 2 1 2 2 2 0 1 1.44
Claude 4 card 05/25 2 1 2 1 2 2 1 0 1 1.33
Gemini 2.5 paper 06/25 1 1 2 1 2 2 1 1 2 1.44
IJTE agent report 07/25 1 1 2 0 2 2 1 0 1 1.11
@@
While all evaluation reports provide at least a general overview of procedures, the level of detail differs. We find that audit reports consistently offer information on evaluation scope (average score 1.38) and procedures (1.29) used. However, contextual details are much less consistently reported (average 0.94, $\sigma=0.68$) both across organizations and sometimes within the same organization. For example, Google DeepMind released four documents with significantly varying levels of contextual detail (average scores of 1.00, 0.67, 0.56, 1.44, and 1.44 in chronological order). Only some reports (e.g., Meinke et al. (2025); Gemma Team et al. (2024); Üstün et al. (2024)) offered detailed information on the scaffolding used to evaluate the system.
Reports inconsistently address assumptions, limitations, and auditor information. While justifications and limitations are more commonly acknowledged (22 of 24) than assumptions (16 of 24), they are typically discussed at a high level rather than relating to specific methodological choices. Ten reports discuss limitations minimally, while twelce do so comprehensively. However, there are notable qualitative differences even among more comprehensive discussions. For example, Üstün et al. (2024) discuss limitations in prompting techniques, language transferability, evaluation tool chaining, and reproducibility in significant detail, whereas the OpenAI system card (OpenAI 2024) only consider limitations of what was being evaluated (e.g. classified information or restricted data) rather than substantial limitations of the procedures. Regarding auditor information, 15 of 24 reports include auditor expertise details, but rarely their background (3 of 24), making an assessment of potential biases in the audit harder.
Reports rarely disclose information about evaluation integrity, resources, review processes, and obsolescence criteria. Only 4 of 24 reports detail organization and auditor selection, conflicts of interest, or contractual arrangements governing the evaluation process. Even in these cases, the report tended to be about the contractual details of human baselines that are part of the evaluation rather than the core auditing team itself. Only 11 of 24 reports specify which resources auditors had access to, including model access level, computational resources, and time constraints. Similarly, only eight reported on some form of review and ten on feedback mechanisms (Meinke et al. 2025; METR 2024; OpenAI 2024; UK AISI and US AISI 2024a). Seven report whether they underwent any form of peer review or quality assurance process before publication (e.g. Gemini Team et al. 2024b; UK AISI and US AISI 2024b). Five specified an option for public feedback and corrections. Zero reports explicitly present obsolescence criteria.
@@ 5 What guidance do current governance frameworks offer?
Principles Features
Document Justif. Assum. Limit. Who What How Access Integrity Review
Regulations
Brazil 0 0 0 0 0 0 0 0 0
China 0 0 0 0 1 1 0 0 1
EU 1 1 1 1 1 1 1 1 1
Singapore 0 0 0 1 1 1 0 0 1
South Korea 0 0 0 0 0 0 0 0 0
USA 0 0 0 0 1 0 0 0 0
New York 0 0 0 0 0 1 0 0 0
Other norms
Bletchley 0 0 0 0 0 1 0 0 1
Int’l Report 1 1 1 1 1 1 1 1 1
Japan AISI 0 0 0 0 1 0 0 0 0
METR 1 1 1 1 1 1 0 1 1
NIST AI RMF 1 1 1 1 1 1 0 1 1
Paris 0 0 0 0 0 0 0 0 0
Seoul 0 0 0 0 1 0 0 0 0
UK AISI 0 0 1 1 1 1 0 0 1
Company policies
Anthropic 1 1 0 0 1 1 0 1 1
Google DeepMind 1 0 1 1 1 1 0 1 1
Meta 1 0 1 1 1 1 0 1 1
Microsoft 1 1 1 1 1 1 0 0 1
OpenAI 1 0 0 0 1 1 0 0 1
xAI 1 0 1 0 1 1 0 0 1
Table 3: What guidance do existing governance frameworks provide for reporting? We give binary scores of 1 (requires or recommends explicit reporting of this feature) or 0 (may mention issue’s importance to the ecosystem but not specifically require or recommend its disclosure or reporting). For company policies we score disclosure of it. Citations and details for the specific documents are in Appendix D.
Methodology: We analyzed key governance frameworks issued by institutions that are considered influential regarding technical evaluations for risks from state-of-the-art models. We selected these frameworks based on jurisdiction, institutional authority, and direct influence on model evaluations (see Appendix D for details). This included official regulations and policies (EU AI Act, US AI Action Plan), voluntary industry standards (NIST AI Risk Management Framework), and policies from major AI developers. We scored each framework on whether it either explicitly required or recommended reporting of each feature in our audit card template, as shown in Table 3.
Existing frameworks offer limited guidance on audit reporting. We find existing frameworks consistently require that models are evaluated (treated synonymously with being assessed or tested) but are not at the level of detail to specifically require or even recommend evaluation reports state the features of the evaluation we have listed. This is especially true for these features: assumptions, limitations, process integrity matters, resources and access given to auditors involved. This is somewhat expected, as governance frameworks for emerging technologies crystallize gradually (Linkov et al. 2018), and the supplementary lower-level details around evaluations in the form of soft law are still emerging.
@@ 6 Challenges with auditing according to expert stakeholders
Methodology: To more thoroughly understand audit reporting and perceived gaps between current and best practices, we interviewed 10 experts. We reached out to these experts based on their expertise and familiarity across the evaluation ecosystem. The 10 experts come from a diverse range of private companies, non-profit organizations, and government bodies. Interviewees include: evaluation designers (P1), evaluation developers (P3, P4, P5), evaluation report writers (P1, P3, P5), evaluation-related policy researchers (P6, P7), and those supporting evaluation delivery and standards development (P8, P9). Interview findings reveal insights that extend beyond that documented in existing literature. To allow for candid discussions, we committed to working with the level of anonymity that was comfortable for experts. See Appendix E for methodological details, interviewees’ backgrounds, interview questions and consent forms.
@@ 7 Discussion and recommendations
There are currently no established standards on how to report the context of audits. The lack of standards leads to inconsistency in audit reporting practices. Certain contextual features, such as reporting on the integrity of auditing processes, are particularly neglected. Lacking transparency around audit methodology and context impedes accountability in the AI ecosystem, where trust in evaluations is essential for informed decision-making by regulators, users, and the public.
Audit cards address AI evaluation reporting gaps to aid methodological rigor and public transparency. Current inconsistent reporting practices undermine evaluations’ governance role, creating trust gaps between developers, auditors, regulators, and the public. By structuring the essential contextual information to be reported, audit cards enable meaningful interpretation of evaluation results. Countries vary in how they see evaluations—emphasizing scientific rigor or compliance verification, public transparency or government oversight. While this paper does not comment on which regulatory approach is superior, it does take the premise that evaluations are scientific exercises requiring public transparency for effective governance.
Audit cards are one approach among several to increase transparency between developers, users, and other stakeholders. Audit cards are designed to offer a structured checklist applicable to all types of assessment (see Appendix C), but should be viewed as an instrument in a larger AI accountability toolkit. The context they offer can be helpful to other mechanisms, including model and system cards, datasheets, and benchmark details (Mitchell et al. 2019; Gebru et al. 2021; Reuel et al. 2024b).
@@
  • Limited sample size and reporting details. To enable consistent high-level analysis of the current ecosystem, we used simple 2 or 3 point scoring methods. However, this comes at the expense of granular details and nuance, differentiating between reports only to a limited degree. This is particularly pronounced for our analysis of governance frameworks, where we use binary scores. Furthermore, our analysis, while extensive (28 academic papers, 20 evaluation reports, and 21 governance frameworks), is not exhaustive.
@@ Acknowledgments
This research was supported by the ML Alignment & Theory Scholars (MATS) Program, which provided funding for Leon Staufer and Mick Yang through research stipends. We also thank MATS and our research managers Juan Gil and Keivan Navaie for their organizational assistance and research support.
@@ Appendix A Disambiguating auditing and the evaluations process
AI audits, for the purpose of our analysis, refer to formalized evaluation processes of AI models or systems. By formalized, we mean that it is performed systematically and by an organization, i.e., not just being performed ad hoc by individuals. This kind of evaluation is a model audit, as opposed to a governance audit or an application audit (Mökander et al. 2024).
The evaluation is the underlying assessment run on the model during the audit. We account for the fact that evaluations can occur internally or externally, and pre- or post-deployment. Evaluations can assess capabilities, propensities, or risks. Evaluations can be in many forms, such as benchmarks, red-teaming, human uplift studies, and user studies. Our approach considers all these variants, as the contextual features we have identified are relevant to all. It is possible, that in some of these contexts certain features will be more important and should therefore be reported in more detail. We explore the importance of certain features further in Section 7 and Appendix C.
@@
When speaking of evaluation reports, we refer to internal or external documents including evaluation results (e.g., a model’s results on a benchmark). Examples of evaluation reports include: (OpenAI 2024; Meinke et al. 2025) as well as private evaluation reports produced by third-parties for AI companies and internal reports created by AI companies themselves.
@@ Appendix B Literature review
Methodology Details
We reviewed past work on “AI auditing”, “AI evaluations”, “best practices”, “science of evaluations”, “evaluation context” and “reporting”. We then assessed each paper for relevance based off the titles. This only yielded eight papers. We thus expanded to adjacent fields such as transparency and AI auditing, technical evaluation design, and sociotechnical approaches. These fields integrate technical performance metrics with societal assessments and considerations of how technical performance interacts with real-world social, cultural, and institutional factors.
We scored recommendations within each paper as major arguments (2), minor arguments (1), and no mention (0). Scoring of (2) represents that the paper actively discussed and/or advocates for a feature, rather than mentioning the feature’s importance in passing (1). We chose our three-point scale to balance reliability and granularity. For all three annotations (papers, reports, and regulations), we deliberated on the appropriate scoring system to ensure it is objective, whilst not giving a false sense of detail.
Through four iterations, we refined which aspects were most salient, focusing particularly on aspects that were actionable for auditors. Our methodology included: (1) two annotators discussing discrepancies before establishing the scoring system, (2) iterating on granularity. We found fine-grained scoring unreliable at higher levels, opting instead for our three-tiered system, and (3) developing detailed scoring rubrics with concrete examples for each category.
Finally, we categorized and prioritized these aspects based on input from experts. The mapping of aspects to features is shown in Appendix B. The final audit card is depicted in Figure 1 and Appendix C.
The guidelines for annotation of audit reports in Section 4 were as follows:
Focus is how well the features are reported not how well they are argued/performed.
(0) no information / not mentioned.
(1) some (even minimal) information. Example: “based on conversations with global experts. […] expert consultants” regarding expertise for biological evaluations in Claude 3 model card (Anthropic 2024, 25).
(2) rich information, meaning you don’t have any important open questions after reading. Example: “red team consists of experts in cybersecurity, […], in addition to multilingual content specialists with backgrounds in integrity issues for specific geographic markets” regarding the expertise feature in Llama 3 research paper (Grattafiori et al. 2024, 48).
This granularity is commensurate with our aim: show from a bird’s-eye view that current aspects of audit cards are inconsistently reported on, not provide a definitive assessment of individual audit reports. We find that awarding one point for a minimal amount of information is not an issue, as we want to be charitable in our assessment of reports while still illustrating a frequent lack of thorough reporting.
List of papers
@@
    Towards Publicly Accountable Frontier LLMs (Anderljung et al. 2023)
@@
    Science of Evals (Apollo Research 2024)
@@
    Declare and Justify: Explicit assumptions in AI evaluations are necessary for effective regulation (Barnett and Thiergart 2024a)
@@
    What AI evaluations for preventing catastrophic risks can and cannot do (Barnett and Thiergart 2024b)
@@
    AI auditing: The Broken Bus on the Road to AI Accountability (Birhane et al. 2024)
@@
    Structured Access for Third-Party Research (Bucknall and Trager 2023)
@@
    Rethink reporting of evaluation results in AI (Burnell et al. 2023)
@@
    Black-Box Access is Insufficient for Rigorous AI Audits (Casper et al. 2024)
@@
    A Survey on Evaluation of Large Language Models (Chang et al. 2023)
@@
    Who Audits the Auditors? Recommendations from a field scan of the algorithmic auditing ecosystem (Costanza-Chock, Raji, and Buolamwini 2022)
@@
    Hard choices in artificial intelligence (Dobbe, Krendl Gilbert, and Mintz 2021)
@@
    Dimensions of Generative AI Evaluation Design (Dow et al. 2024)
@@
    Can We Trust AI Benchmarks? (Eriksson et al. 2025)
@@
    The TRIPOD-LLM Reporting Guideline for Studies Using Large Language Models (Gallifant et al. 2025)
@@
    Datasheets for Datasets (Gebru et al. 2021)
@@
    Devising ML Metrics (Hendrycks and Woodside 2024)
@@
    Responsible Reporting for Frontier AI Development (Kolt et al. 2024)
@@
    Holistic Evaluation of Language Models (Liang et al. 2023)
@@
    Task Development Guide (METR b)
@@
    Auditing Large Language Models: A Three-Layered Approach (Mökander et al. 2024)
@@
    Reasons to Doubt the Impact of AI Risk Evaluations (Mukobi 2024)
@@
    Towards AI Accountability Infrastructure (Ojewale et al. 2024)
@@
    Closing the AI Accountability Gap: Defining an End-to-End Framework for Internal Algorithmic Auditing (Raji et al. 2020)
@@
    BetterBench (Reuel et al. 2024b)
@@
    Fairness and Abstraction in Sociotechnical Systems (Selbst et al. 2019)
  26. 26.
    Model evaluation for extreme risks (Shevlane et al. 2023)
  27. 27.
    Toward an evaluation science for generative AI systems (Weidinger et al. 2025)
  28. 28.
    Evaluatology: The Science and Engineering of Evaluation (Zhan et al. 2024)
Paper 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
Anderljung et al.2023 2 0 0 2 2 2 2 2 2 2 2 2 0 2 0 0 2 2 0 0 2 1 2 2
Apollo Research2024 0 0 0 0 0 2 0 0 0 0 2 0 0 0 0 0 0 0 0 0 2 2 0 0
Barnett and Thiergart2024a 0 0 0 0 1 0 0 0 0 0 0 1 0 0 2 0 0 0 0 0 1 2 0 0
Barnett and Thiergart2024b 2 2 0 1 0 0 0 0 0 0 2 0 0 0 0 2 2 0 0 0 2 1 2 0
Birhane et al.2024 1 2 0 0 1 2 0 0 0 0 0 2 0 0 2 0 2 2 0 0 2 2 2 2
Bucknall and Trager2023 2 2 0 2 0 2 2 0 0 0 0 0 2 0 0 0 2 0 0 0 2 0 2 2
Burnell et al.2023 0 0 0 0 0 1 2 0 0 1 2 0 0 2 2 2 0 0 1 0 2 2 0 1
Casper et al.2024 2 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 2 1 0 1
Chang et al.2023 0 0 0 0 1 0 0 0 2 0 0 2 0 0 2 0 0 0 0 0 1 0 0 0
Costanza-Chock, Raji, and Buolamwini2022 2 1 0 0 2 1 0 1 0 0 0 1 0 0 0 0 0 2 0 0 2 2 0 2
Dobbe, Krendl Gilbert, and Mintz2021 0 0 0 0 0 0 0 0 0 0 2 0 0 2 2 0 2 2 1 2 2 2 1 2
Dow et al.2024 0 0 1 0 1 2 0 0 2 0 1 1 0 0 0 0 0 1 0 0 2 0 2 0
Eriksson et al.2025 0 1 0 0 1 2 0 0 1 0 2 1 0 1 2 2 1 0 2 2 2 2 0 0
Gallifant et al.2025 0 0 0 1 2 2 2 2 1 0 2 1 0 2 1 1 2 0 1 0 2 2 2 2
Gebru et al.2021 2 1 1 0 1 1 2 2 2 1 0 1 2 2 2 2 2 2 0 0 1 1 0 2
Hendrycks and Woodside2024 0 0 0 0 0 0 0 0 1 1 2 0 0 0 0 0 1 0 0 0 0 0 0 1
Kolt et al.2024 0 2 0 0 2 0 2 0 0 0 0 0 0 0 0 0 0 0 1 0 1 1 1 2
Liang et al.2023 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 2 0 2 1 0 0
METRb 0 0 0 0 2 2 2 0 1 2 2 2 2 0 0 0 0 0 0 0 0 0 0 0
Mökander et al.2024 2 2 0 0 1 2 2 0 0 0 1 1 0 0 0 0 2 1 2 0 2 0 2 2
Mukobi2024 2 2 2 2 0 2 0 2 0 0 2 0 0 1 0 0 0 0 1 0 2 0 2 1
Ojewale et al.2024 2 1 1 0 1 1 1 0 1 1 1 1 0 0 0 0 0 2 1 0 2 1 0 1
Raji et al.2020 1 0 0 0 0 2 0 0 2 0 2 1 0 0 2 0 2 2 2 2 2 0 0 1
Reuel et al.2024b 0 0 0 0 2 0 2 0 0 2 2 2 2 2 2 0 2 0 0 0 1 0 2 2
Selbst et al.2019 0 0 0 0 0 0 0 0 2 0 2 1 0 2 2 0 2 2 0 2 2 0 0 0
Shevlane et al.2023 2 2 0 0 2 2 0 0 0 0 2 0 0 0 0 1 2 0 0 0 2 0 2 2
Weidinger et al.2025 0 0 0 0 0 2 1 0 0 0 2 0 0 2 0 0 2 0 2 1 2 2 1 0
Zhan et al.2024 0 0 0 0 1 2 2 0 0 0 2 0 0 2 2 2 2 0 0 0 2 2 0 0
Table 4: Annotations for all 28 papers across the 24 initial contextual aspects (see Appendix B). We score each paper’s discussion of these as 2 (major argument), 1 (minor argument), and 0 (no mention).
Mapping of initial aspects
Our initial literature review yielded 24 aspects listed below that were mapped ($\rightarrow{}$) to the principles and features: